OpenStudy (anonymous):

2[a(a + b) + b(a + b)]

OpenStudy (anonymous):

can anyone help me

OpenStudy (anonymous):

yes tell me, do u want to simplify this equation ? If yes then it comes out to be \[2(a+b)^2\] It can be also written as \[2a^2+4ab+2b^2\]

OpenStudy (anonymous):

can u show me in steps please

OpenStudy (anonymous):

where did u get the 4 from

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

We have, 2[a(a + b) + b(a + b)] Let 2 be outside the square brackets and we will solve the part inside the square brackets, \[2[a^2+ab+ab+b^2]\] Now, solving further \[=2 [ a^2 + 2ab + b^2 ]\] \[=2(a+b)^2\] (because we are aware about this algebraic identity \[(a+b)^2=a^2+2ab+b^2\] )

OpenStudy (anonymous):

2[a(a + b) + b(a + b)] from here you can factor out a+b, this becomes =2(a+b)[a+b] and just like Sid has said it became =2(a+b)^2 di you see it A candi?

OpenStudy (anonymous):

so what would the answer b

OpenStudy (anonymous):

is there any answer selection?

OpenStudy (anonymous):

no solve till i cant brake it dwn anymore

OpenStudy (anonymous):

wher does the 4 come from

OpenStudy (anonymous):

Ok I'll tell you, When you solve this \[2(a^2+ab+ab+b^2)\]\[=2(a^2+2ab+b^2)\]\[=2(a^2)+2(2ab)+2(b^2)\]\\[=2a^2+4ab+2b^2\] Is it clear now that where did 4 come from ?

OpenStudy (anonymous):

the 4 came from 2(2ab)=4ab

OpenStudy (anonymous):

Yes you are right